OpenStudy (anonymous):

The sum of the squares of two consecutive positive even integers is 340. Write an equation and find the integers.

OpenStudy (anonymous):

\[(2n)^2+(2n+2)^2=340\]

OpenStudy (anonymous):

\[8n^2+8n+4=340\] \[8n^2+8n-366=0\] \[n^2+n-45=0\] \[(n-6)(n+7)=0\] \[n=6\] or \[n=-7\] choose n = 6